# Towards Black-Box Membership Inference Attack for Diffusion Models

### v3

@@ 5 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment setup in previous papers [6, 20]. The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
5.1 Evaluation Metrics
We follow the metrics used in previous papers [6, 20], including Area Under Receiver Operating Characteristic (AUC), Attack Success Rate (ASR), the True Positive Rate (TP) when the False Positive Rate is 1%. We also plot the ROC curves.
5.2 MIA with DDIM Models
We follow the experimental setup of [6, 20]. We train a DDIM model on the CIFAR-10/100 [22] and STL10-Unlabeled datasets [2], using the image generation step $T=1000$ and sampling interval $k=100$. For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members. The remaining 50% are utilized as nonmembers. We use [30], [6], [20] as our baseline methods. We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$. We will discuss the impact of the diffusion step and the average number in Section 5.5.
Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method CIFAR10 CIFAR100 STL10
Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
Loss [30] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
SecMI [6] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
PIA [20] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
PIAN [20] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
For the difference function $D(x,\hat{x})$, following the setup in [6], we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network [11] and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
5.3 MIA with Diffusion Transformers
We train a diffusion transformer model on the ImageNet [3] dataset following the setup of [33]. We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$. For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
We use [30], [6], [20] as our baseline methods. Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation. For the difference function $D(x,\hat{x})$, following the setup in [6, 20], we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
@@ 5.3 MIA with Diffusion Transformers
Loss [30] $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
SecMI [6] $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
PIA [20] $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
PIAN [20] $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
@@
We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet $256\times 256$ in Figure 3. The curves further demonstrate the effectiveness of our method.
5.4 MIA with the Stable Diffusion Model
We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experiment setup of [6, 20], use the LAION-5B dataset [40] as member and COCO2017-val [25] as non-member. We randomly select 2500 images in each dataset. We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP [23], which we denote as Laion5 with BLIP.
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [48] to measure the differences between two images. The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net.
@@ 5.4 MIA with the Stable Diffusion Model
Loss [30] $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
SecMI [6] $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
PIA [20] $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
PIAN [20] $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
@@ 5.5 Ablation Studies
Figure 4: The impact of average numbers. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer model on Imagenet. In both cases, averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
@@
Figure 6: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Stable Diffusion model on LAION-5. We find that adjusting the sampling interval has a relatively small influence in the first case and does not affect our method in the latter case. This makes our algorithm applicable to more setups.
@@ 6 An Application to DALL-E’s API
In this section, we conduct a small experiment with online API services to test the effectiveness of our algorithm. We test with the DALL-E 2 [35] model since DALL-E 2 provides a variation API service.
We select different thresholds and classify an image with a variation error below the threshold as members and those above the threshold as non-members. We then calculate metrics such as AUC and ASR. Since the baseline algorithms [30, 6, 20] require intermediate results, we are unable to test these algorithms under the online API setup.
One challenge with MIA for DALL-E 2 is that it does not disclose its training set. However, since it is adept at generating various famous artworks, we select 30 famous paintings from five famous artists: Van Gogh, Monet, Da Vinci, Dali, and Rembrandt, to form our member set. We believe that it is reasonable to hypothesize that these artworks are used in DALL-E 2’s training set. For constructing the non-member set, we used Stable Diffusion 3 [9] to generate images based on the titles of each painting in the member set. The benefit of constructing non-members in this way is that it allows for control over the content of the artwork descriptions, reducing bias caused by content shift. Moreover, these generated images are certainly not in the DALL-E 2’s training set. Other results of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
The results, presented in Table 4, demonstrate that our algorithm achieves a relatively high accuracy under this evaluation method. Our observation is illustrated in Figure 17. As seen in the figure, for Monet’s iconic painting "Water Lily Pond", the original artwork shows minimal changes when using DALL-E 2’s variation API, retaining most of its main features. In contrast, the artwork generated by Stable Diffusion 3 undergoes significant changes, with variations in both the number of flowers and lily pads. Therefore, we hypothesize that artworks with smaller changes after API usage are more likely to have appeared in the model’s training set.
Table 4: The results of applying our algorithm to DALL-E 2’s variation API. We assume some famous paintings as members and use Stable Diffusion 3 along with the titles of these artworks to generate corresponding non-members. Our algorithm also achieves high accuracy under this setup. Since the baseline algorithm requires noise prediction results, we are unable to evaluate it in this black-box setup.
@@
Figure 7: Main observation of our attack. DALL-E 2’s variation API makes minimal changes to famous artworks, while nonmember images with similar content undergo significant alterations.
@@ Appendix A Datasets, Models and Hyperparameters
For DDIM, we follow the training hyperparameters of [6] to train a DDIM model on the CIFAR-10/100 [22] and STL10-Unlabeled [2] datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$. The training iterations are set to 800,000. For all the datasets, we randomly select 50% of the training samples to train the model and designate them as members. The remaining 50% of the training samples are used as nonmembers.
We use [30, 6, 20] as our baseline methods. We use their official code repositories and apply the optimal hyperparameters from their papers. For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to average the output images as $\hat{x}$.
For the Diffusion Transformers, we train the model on the ImageNet [3] dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of [33]. We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$. For the $128\times 128$ image size, we use 160,000 training iterations. For the $256\times 256$ image size, we use 300,000 training iterations. These numbers of training iterations are chosen to ensure the generation of high-quality images.
@@
We also use [30, 6, 20] as our baseline methods. Since these papers do not study Diffusion Transformers, we adapt their algorithms to the DiT framework for evaluation. We fix the DDIM step at $k=50$ and choose diffusion steps $t\in[50,100,150,200,250,300]$, recording their optimal solutions under these different hyperparameter settings.
For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experimental setup of [6, 20], using an image generation step of $T=1000$ and a sampling interval of $k=10$. We use the LAION-5B dataset [40] as the member set and COCO2017-val [25] as the non-member set. We randomly select 2500 images from each dataset. We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP [23], denoted as Laion with BLIP.
We use the hyperparameters from the papers [6, 20] to run the baseline methods. For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [48] to measure the differences between two images. Other hyperparameters remain the same as in the baseline methods.
@@ Appendix B More Experiment Results
Besides the ROC curves in the main paper, we also plot ROC curves for the Diffusion Transformer train on ImageNet $128\times 128$ and the Stable Diffusion train on Laion5 in Figure 16. The curves further demonstrate the effectiveness of our method.
Figure 16: The ROC curves of various setups. Left: Diffusion Transformer on ImageNet $128\times 128$. Right: The Stable Diffusion on Laion5. The curves show that our algorithm outperforms the baseline algorithms.
@@ Appendix C Proof
In this section, we implement the proof of the theorem. See 1
@@
If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$. From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV [8], we get the following probability inequality for any $\beta>0$:

### v4

@@ 5 Experiments
Loss (Matsumoto et al., 2023) $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
SecMI (Duan et al., 2023) $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
PIA (Kong et al., 2023) $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
PIAN (Kong et al., 2023) $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
@@
Loss (Matsumoto et al., 2023) $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
SecMI (Duan et al., 2023) $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
PIA (Kong et al., 2023) $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
PIAN (Kong et al., 2023) $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
@@
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment setup in previous papers (Duan et al., 2023; Kong et al., 2023). The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
5.1 Evaluation Metrics
We follow the metrics used in previous papers (Duan et al., 2023; Kong et al., 2023), including Area Under Receiver Operating Characteristic (AUC), Attack Success Rate (ASR), the True Positive Rate (TP) when the False Positive Rate is 1%. We also plot the ROC curves.
5.2 MIA with DDIM Models
We follow the experimental setup of (Duan et al., 2023; Kong et al., 2023). We train a DDIM model on the CIFAR-10/100 (Krizhevsky et al., 2009) and STL10-Unlabeled datasets (Coates et al., 2011), using the image generation step $T=1000$ and sampling interval $k=100$. For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members. The remaining 50% are utilized as nonmembers. We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods. We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$. We will discuss the impact of the diffusion step and the average number in Section 5.5.
For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network (He et al., 2016) and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
5.3 MIA with Diffusion Transformers
We train a diffusion transformer model on the ImageNet (Deng et al., 2009) dataset following the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$. For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods. Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation. For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet in Appendix B. The curves further demonstrate the effectiveness of our method.
5.4 MIA with the Stable Diffusion Model
We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experiment setup of (Duan et al., 2023; Kong et al., 2023), use the LAION-5B dataset (Schuhmann et al., 2022) as member and COCO2017-val (Lin et al., 2014) as non-member. We randomly select 2500 images in each dataset. We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP (Li et al., 2022), which we denote as Laion5 with BLIP.
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images. The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net.
@@ 5.5 Ablation Studies
Figure 3: The impact of average numbers. We train DDIM model on CIFAR-10. Averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
@@
Figure 5: The impact of sampling intervals. We train DDIM model on CIFAR-10. We find that adjusting the sampling interval has a relatively small influence in the first case and does not affect our method in the latter case. This makes our algorithm applicable to more setups.
@@ 6 An Application to DALL-E’s API
In this section, we conduct a small experiment with online API services to test the effectiveness of our algorithm. We test with the DALL-E 2 (Ramesh et al., 2022) model since DALL-E 2 provides a variation API service. We select different thresholds and classify an image with a variation error below the threshold as members and those above the threshold as non-members. We then calculate metrics such as AUC and ASR. Since the baseline algorithms (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) require intermediate results, we are unable to test these algorithms under the online API setup.
One challenge with MIA for DALL-E 2 is that it does not disclose its training set. However, since it is adept at generating various famous artworks, we select 30 famous paintings from five famous artists: Van Gogh, Monet, Da Vinci, Dali, and Rembrandt, to form our member set. We believe that it is reasonable to hypothesize that these artworks are used in DALL-E 2’s training set. For constructing the non-member set, we used Stable Diffusion 3 (Esser et al., 2024) to generate images based on the titles of each painting in the member set. The benefit of constructing non-members in this way is that it allows for control over the content of the artwork descriptions, reducing bias caused by content shift. Moreover, these generated images are certainly not in the DALL-E 2’s training set.
The results, presented in Table 4, demonstrate that our algorithm achieves a relatively high accuracy under this evaluation method. Our observation is illustrated in Figure 6. As seen in the figure, for Monet’s iconic painting ”Water Lily Pond”, the original artwork shows minimal changes when using DALL-E 2’s variation API, retaining most of its main features. In contrast, the artwork generated by Stable Diffusion 3 undergoes significant changes, with variations in both the number of flowers and lily pads. Therefore, we hypothesize that artworks with smaller changes after API usage are more likely to have appeared in the model’s training set. Other results of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
Table 4: The results of applying our algorithm to DALL-E 2’s variation API. We assume some famous paintings as members and use Stable Diffusion 3 along with the titles of these artworks to generate corresponding non-members. Our algorithm also achieves high accuracy under this setup.
@@
Figure 6: Main observation of our attack. DALL-E 2’s variation API makes minimal changes to famous artworks, while nonmember images with similar content undergo significant alterations. More examples of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
@@ 7 Conclusion, Limitations and Future Directions
Impact Statement
This paper addresses the critical issue of protecting intellectual property and data privacy in the era of rapidly advancing generative models. By proposing a novel membership inference attack (MIA) method, our work aims to identify whether specific artworks have been used during the training of diffusion models. This has significant implications for safeguarding creators’ copyrights, detecting unauthorized use of artistic works, and promoting ethical practices in the development and deployment of generative models. While our method is designed with privacy protection in mind, we acknowledge the potential for misuse. We hope this research serves as a foundation for further discussions on data privacy and ethical considerations, encouraging responsible use and development of diffusion models in the community.
@@ Appendix A Datasets, Models and Hyperparameters
For DDIM, we follow the training hyperparameters of (Duan et al., 2023) to train a DDIM model on the CIFAR-10/100 (Krizhevsky et al., 2009) and STL10-Unlabeled (Coates et al., 2011) datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$. The training iterations are set to 800,000. For all the datasets, we randomly select 50% of the training samples to train the model and designate them as members. The remaining 50% of the training samples are used as nonmembers.
We use (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) as our baseline methods. We use their official code repositories and apply the optimal hyperparameters from their papers. For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to average the output images as $\hat{x}$.
For the Diffusion Transformers, we train the model on the ImageNet (Deng et al., 2009) dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$. For the $128\times 128$ image size, we use 160,000 training iterations. For the $256\times 256$ image size, we use 300,000 training iterations. These numbers of training iterations are chosen to ensure the generation of high-quality images.
@@
We also use (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) as our baseline methods. Since these papers do not study Diffusion Transformers, we adapt their algorithms to the DiT framework for evaluation. We fix the DDIM step at $k=50$ and choose diffusion steps $t\in[50,100,150,200,250,300]$, recording their optimal solutions under these different hyperparameter settings.
For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experimental setup of (Duan et al., 2023; Kong et al., 2023), using an image generation step of $T=1000$ and a sampling interval of $k=10$. We use the LAION-5B dataset (Schuhmann et al., 2022) as the member set and COCO2017-val (Lin et al., 2014) as the non-member set. We randomly select 2500 images from each dataset. We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP (Li et al., 2022), denoted as Laion with BLIP.
We use the hyperparameters from the papers (Duan et al., 2023; Kong et al., 2023) to run the baseline methods. For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images. Other hyperparameters remain the same as in the baseline methods.
@@ Appendix B More Experiment Results
We plot ROC curves for the DDIM train on CIFAR-10 and Diffusion Transformer train on ImageNet $256\times 256$ in Figure 15. The curves further demonstrate the effectiveness of our method.
@@ Appendix C Proof
In this section, we implement the proof of the theorem. See 4.2
@@
If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$. From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV (Durret, 2010), we get the following probability inequality for any $\beta>0$: